A pitot static tube is used to measure the velocity of air flowing through a duct. The manometer shows a difference in head of 5 cm of water. If density of air and water are 1.13 kg/m3 and 1000

kg/m3, determine the velocity of air. Assume the coefficient of discharge of the Pitot tube as 0.98.


1
Expert's answer
2021-07-20T09:40:39-0400

h=5cm

"\rho_w=1000kg/m^3\\\rho_a=1.13kg/m^3\\C_v=0.98"

We know that

"V=C_v\sqrt{2gh(\frac{\rho_w}{\rho_a}-1})"

Put value


"V=0.98\times\sqrt{2\times9.8\times0.05\times(\frac{1000}{1.13}-1)}=28.86m/sec"